# Model of incompressible turbulent flows via a kinetic theory

### v1

@@ 5.3 Non-equilibrium characteristics and velocity distribution functions
The non-equilibrium and non-Newtonian characteristics predicted by the present kinetic model can be essentially explained from the VDFs. We now examine the reduced VDFs obtained from both the LR-BGK and HR-BGK models at $\text{Re}=10133$, comparing the computed VDFs with their corresponding local equilibrium distributions to quantify the degree of departure from equilibrium.
@@
Figures 8 and 9 present the reduced velocity distribution functions $\Phi_{1}$, $\Phi_{2}$, and $\Phi_{3}$ as functions of microscopic velocity $\xi_{y}$ in various normal wall positions ranging from the viscous sublayer to the centerline of the channel, compared against the local equilibrium VDFs derived from the local mean velocity and turbulent kinetic energy. Here $\xi_{y}$ in the region $\left[-3\xi_{m},3\xi_{m}\right]$ with $\xi_{m}=\sqrt{4K/3}$ represents the most probable velocity scale based on the local turbulent kinetic energy. The fourth VDF $\Phi_{4}$, which exhibits behaviour similar to $\Phi_{1}$ by remaining close to equilibrium, is presented in Appendix C.
@@
The most striking observation is the markedly different behaviours exhibited by the various reduced VDFs. The distribution function $\Phi_{1}$, which represents the zeroth-order moment in the wall-normal direction, remains remarkably close to its local equilibrium state across all wall-normal positions and for both kinetic models. This near-equilibrium behaviour indicates that the mass conservation constraint is well satisfied throughout the flow domain, consistent with the incompressibility condition. In sharp contrast, $\Phi_{2}$ and $\Phi_{3}$ exhibit pronounced departures from their respective equilibrium distributions, particularly in the outer region of the flow. The VDF $\Phi_{2}$, associated with the streamwise momentum transport, shows increasing non-equilibrium effects as the distance from the wall increases. Notably, at $y^{+}\approx 419$ and the channel centerline, $\Phi_{2}$ exhibits distinctly non-Gaussian distributions, indicating the presence of intermittent large-scale turbulent structures that cannot be captured by simple equilibrium assumptions. Similarly, $\Phi_{3}$, which is related to the streamwise kinetic energy transport, demonstrates substantial non-equilibrium characteristics with a comparable spatial evolution pattern. The VDF $\Phi_{3}$ shows progressively stronger departures from equilibrium as the wall-normal distance increases, with the outer region exhibiting the most pronounced non-equilibrium effects. Remarkably, $\Phi_{3}$ also displays non-Gaussian distributions at $y^{+}\approx 419$ and the channel centerline. The non-equilibrium nature of $\Phi_{3}$ directly contributes to the anisotropic stress behaviour observed in the Reynolds stress predictions, explaining the model’s capability to capture the non-Newtonian effects that linear eddy viscosity models cannot represent.
The comparison between the LR-BGK and HR-BGK models reveals consistent non-equilibrium patterns, with the LR-BGK model providing slightly more detailed resolution of the near-wall non-equilibrium effects due to its finer grid resolution in the viscous sublayer. Overall, both kinetic models successfully capture the essential physics of the non-equilibrium turbulent transport processes, demonstrating the robustness of the kinetic approach across different near-wall treatments and grid resolutions. To further quantify the degree of departure from equilibrium, we examine the second-order moment of $F-F^{eq}$
@@
Figures 10 and 11 present the deviatoric part of $\bm{\sigma}^{neq}$ predicted by LR-BGK model at $\text{Re}=10133$ and compare them with the CE expansion results given by equation (26). It can be seen that the deviatoric Reynolds stress shows good agreement between the LR-BGK model and the CE expansion throughout the flow domain, validating the theoretical consistency of the kinetic approach. The minor discrepancies observed can be attributed to higher-order terms in the CE expansion that are not captured by the second-order analysis. Furthermore, the deviatoric Reynolds normal stress components better reflect the degree of departure from non-equilibrium of the VDFs mentioned above. In the viscous sublayer and buffer region, the system remains essentially in equilibrium with all three normal stress components close to zero, while in the outer layer non-equilibrium effects become more pronounced as the deviations from zero increase progressively.
These non-equilibrium characteristics, as revealed through both the VDFs and their non-equilibrium moments, provide a fundamental explanation for the superior performance of the kinetic models in capturing the anisotropic Reynolds stress behaviour and the associated non-Newtonian effects. The departure from local equilibrium enables the kinetic approach to naturally account for the complex momentum and energy transport mechanisms that arise from the finite relaxation times of turbulent eddies, analogous to the finite collision times in rarefied gas flows. This non-equilibrium characteristic is particularly important in wall-bounded flows, where the presence of the wall creates strong gradients and anisotropic conditions that drive the system away from local Gaussian equilibrium.
Figure 8: The reduced velocity distribution functions $\Phi_{1}$ (first column), $\Phi_{2}$ (second column) and $\Phi_{3}$ (third column) from the LR-BGK model at $\text{Re}=10133$ as a function of the microscopic velocity $\xi_{y}$ at different wall-normal positions from top to bottom: $y^{+}=0.48$ (viscous sublayer), $y^{+}=11.6$ (buffer layer), $y^{+}=215.5$ (log-law region), $y^{+}=419.8$ (log-law region), and $y^{+}=460.4$ (channel centerline). Dash lines represent the computed distribution functions, while solid lines represent the local equilibrium distribution functions.
@@
Figure 9: The reduced velocity distribution functions $\Phi_{1}$ (first column), $\Phi_{2}$ (second column) and $\Phi_{3}$ (third column) from the HR-BGK model at $\text{Re}=10133$ as a function of the microscopic velocity $\xi_{y}$ at different wall-normal positions from top to bottom: $y^{+}=11.3$ (buffer layer), $y^{+}=215.3$ (log-law region), $y^{+}=419.2$ (log-law region), and $y^{+}=453.2$ (channel centerline). Dash-dot lines represent the computed distribution functions, while solid lines represent the local equilibrium distribution functions.
@@
Figure 10: Deviatoric Reynolds shear stress component in wall units of turbulent Couette flow at $\text{Re}=10133$. The solid lines with open circles and dash-dot lines represent the predictions of the LR-BGK model (59) and the CE expansion results (60), respectively. Here, the deviatoric Reynolds stress is defined as $\bm{\sigma}^{d}=\bm{\sigma}-\frac{1}{3}\left(\text{tr}\bm{\sigma}\right)\bm{I}$.
Figure 11: Deviatoric Reynolds normal stress components in wall units of turbulent Couette flow at $\text{Re}=10133$. The solid lines with circles, squares, and triangles represent ${\sigma}_{xx}^{d}$, ${\sigma}_{yy}^{d}$, and ${\sigma}_{zz}^{d}$ predicted by the LR-BGK model (59), respectively. The dash-dot, dash and dot lines represent ${\sigma}_{xx}^{d}$, ${\sigma}_{yy}^{d}$, and ${\sigma}_{zz}^{d}$ predicted by the CE expansion results (60), respectively. Here, the deviatoric Reynolds stress is defined as $\bm{\sigma}^{d}=\bm{\sigma}-\frac{1}{3}\left(\text{tr}\bm{\sigma}\right)\bm{I}$.
@@ 6 Conclusions
In this work, we extended and analyzed the kinetic model for incompressible turbulence proposed by chen2023average and demonstrated improved applicability and physical consistency. By redefining the relaxation time as $\tau=K/(7\epsilon)$, the Chapman–Enskog expansion of the kinetic model yielded linear and nonlinear eddy viscosity and gradient diffusion closures with transport coefficients ($C_{\mu}=0.0816,\text{Pr}_{T}=0.7$) that are in quantitative agreement with those of conventional turbulence models. The analysis of our kinetic model reveals that its second-order expansion generates various material derivatives and higher-order derivatives of TKE. Unlike previous results for the Boltzmann-BGK equation without TKE dissipation (chen2004expanded) or for the Fokker-Planck equation (heinz2007unified; luan2025constructing), the analysis of our kinetic model yielded different forms of material derivatives for the linear terms and higher-order derivative terms involving TKE from the second-order expansion. To our knowledge, such closed-form Chapman–Enskog expressions for the TKE flux and its nonlinear corrections have not previously been reported for this class of turbulent kinetic models.
To address wall-bounded turbulence, we developed two approaches within a unified kinetic framework. For applications outside the viscous sublayer, the non-equilibrium extrapolation boundary condition based on wall function approaches was proposed for the HR-BGK model, maintaining accurate logarithmic layer behaviour at lower computational expense. For direct viscous sublayer resolution, we constructed a low-Reynolds-number kinetic model incorporating a damping function and a viscous diffusion term, with the fully diffuse wall condition. It is worth noting that the present methods for wall-bounded flows are fundamentally different from the boundary conditions based on the wall function method of srinivasan1977turbulent, devised for the kinetic model of lundgren1969model. Our approach is not restricted to regions outside the viscous sublayer, enabling treatment of the complete flow domain from wall to bulk region.
@@
The present kinetic models were comprehensively validated against experimental data and DNS data for turbulent plane Couette flow, demonstrating accurate prediction of mean velocity profiles, skin friction coefficients, and Reynolds stress distributions. They can capture essential flow physics including logarithmic law behaviour and Reynolds stress anisotropy. Although not as pronounced as indicated by DNS, the present kinetic theoretic models are shown to predict the separations between the normal stress components. In terms of predictive performance, the kinetic models achieve comparable accuracy to conventional nonlinear eddy viscosity models  (speziale1987nonlinear; nisizima1987turbulent; myong1990prediction; craft1996development) for mean velocity profiles, skin friction coefficients, and Reynolds shear stress components. Nevertheless, accurate prediction of individual normal stress components in canonical wall-bounded flows remains a shared challenge. A distinctive advantage of the kinetic approach lies in its robust theoretical foundation, requiring only a single relaxation time, whereas nonlinear eddy viscosity models necessitate multiple adhocad\ hoc empirical coefficients for closure. These findings suggest that future developments should explore higher-order information about the collision term $\mathcal{C}(F)\equiv-\nabla_{\bm{\xi}}\cdot\langle\bm{a}^{\prime}f^{\prime}\rangle$ to achieve more accurate modelling of stress tensor anisotropy.
@@
Analysis of the reduced velocity distribution functions revealed significant departures from local Gaussian equilibrium distribution, particularly for streamwise momentum and energy transport components. The non-equilibrium moments obtained from the kinetic model show excellent quantitative agreement with second-order Chapman-Enskog predictions, confirming the theoretical self-consistency of the approach. These non-equilibrium characteristics explain the ability of the kinetic model to capture non-Newtonian effects that linear eddy viscosity models cannot represent, highlighting the fundamental advantage of the kinetic approach over conventional turbulence modelling strategies.
@@
Although this study focused on simple shear flows, the kinetic framework provides a foundation for modelling more complex turbulent phenomena. Future work should address three-dimensional flows, unsteady effects, and extension to compressible turbulence, fully exploiting the theoretical advantages of the kinetic approach. The kinetic turbulence model represents a significant step toward physics-based turbulence modelling, offering reduced dependence on empirical parameters while maintaining computational tractability for engineering applications. We wish to emphasize that the present work is to propose a kinetic theory based turbulence model, in the form of a Boltzmann-BGK differential equation, for better and more naturally capturing certain fundamental underlying physics of the averaged dynamics of turbulence, instead of a kinetic theory based computational method for numerically solving any turbulence problems
@@ Appendix A Derivations of the Reynolds stress and turbulent kinetic energy flux via Chapman–Enskog expansion
$Q^{eq}_{\alpha}$ $=\frac{1}{2}\int\left(\xi_{\alpha}-U_{\alpha}\right)\left(\xi_{\beta}-U_{\beta}\right)^{2}F^{eq}d\bm{\xi}=0$
@@
$\frac{\partial F^{eq}}{\partial U_{\beta}}$ $=\frac{3\left(\xi_{\beta}-U_{\beta}\right)}{2K^{eq}}F^{eq},$ (72)
$\frac{\partial F^{eq}}{\partial K^{eq}}$ $=\frac{3}{2K^{eq}}\left(\frac{\left(\xi_{\alpha}-U_{\alpha}\right)^{2}}{2K^{eq}}-1\right)F^{eq},$
@@
$F^{(1)}=$ $-\tau\left(\partial_{t_{1}}+\xi_{\alpha}\frac{\partial}{\partial x_{1\alpha}}+\bar{a}_{\alpha}^{(1)}\frac{\partial}{\partial\xi_{\alpha}}\right)F^{eq}$ (73)
$=$ $-\frac{3\tau\left(\xi_{\alpha}-U_{\alpha}\right)}{2K^{eq}}\left(\frac{\left(\xi_{\alpha}-U_{\alpha}\right)^{2}}{2K^{eq}}-\frac{5}{3}\right)\frac{\partial K^{eq}}{\partial x_{1\alpha}}F^{eq}$
@@
$\sigma^{(1)}_{\alpha\beta}$ $=-\frac{2}{3}K^{(1)}\delta_{\alpha\beta}+\frac{4\tau K^{eq}}{3}S_{\alpha\beta}^{(1)},$ (74)
$Q_{\alpha}^{(1)}$ $=-\frac{10}{9}\tau K^{eq}\frac{\partial{K^{eq}}}{\partial x_{1\alpha}}.$
@@
$\Phi_{1}\left(y,\xi_{y},t\right)$ $=\int F\left(\bm{x},\bm{\xi},t\right)d\xi_{x}d\xi_{z},$ (81)
$\Phi_{2}\left(y,\xi_{y},t\right)$ $=\int\xi_{x}F\left(\bm{x},\bm{\xi},t\right)d\xi_{x}d\xi_{z},$
$\Phi_{3}\left(y,\xi_{y},t\right)$ $=\int\xi_{x}^{2}F\left(\bm{x},\bm{\xi},t\right)d\xi_{x}d\xi_{z},$
$\Phi_{4}\left(y,\xi_{y},t\right)$ $=\int\xi_{z}^{2}F\left(\bm{x},\bm{\xi},t\right)d\xi_{x}d\xi_{z}.$
@@
$\sigma_{xx}\left(y,t\right)$ $=-\int\left(\Phi_{3}-U_{x}^{2}\Phi_{1}\right)d\xi_{y}+2U_{x}^{2},$ (83)
$\sigma_{xy}\left(y,t\right)$ $=-\int\xi_{y}\Phi_{2}d\xi_{y}+U_{x}U_{y},$
$\sigma_{yy}\left(y,t\right)$ $=-\int\left(\xi_{y}-U_{y}\right)^{2}\Phi_{1}d\xi_{y},$
$\sigma_{zz}\left(y,t\right)$ $=-\int\Phi_{4}d\xi_{y}.$
@@
where $\Delta y=y_{j+1/2}-y_{j-1/2}$ is the length of the cell $j$, $\Phi_{j-1/2}^{n+1/2}$ and $\Phi_{j+1/2}^{n+1/2}$ are the left and right VDFs at the cell interface at time $t^{n+1/2}$, respectively.
To eliminate the implicit collision term, we introduce the transformed VDFs
$\tilde{\Phi}_{\alpha}=\Phi_{\alpha}-\frac{\Delta t}{\tau}\left(\Phi^{eq}_{\alpha}-\Phi_{\alpha}\right),$ (88)
and the mean variables can be computed by taking moments of the transformed VDFs,
@@
As a result, the original VDFs and the mean variables at $t^{n+1}$ can be obtained from $\tilde{\Phi}^{n+1}_{\alpha,j}$ according to (88) and (89),
@@
With the Taylor expansion around the cell interface $y_{j+1/2}$, for smooth flows, $\Phi_{\alpha}\left(y_{j+1/2},\xi_{y},t^{n+1/2}\right)$ can be approximated by linear interpolations,
@@
Then, the mean variables and $\epsilon$ at the cell interface are also obtained by linear interpolation,
$\bm{U}_{j+1/2}^{n}=\bm{U}_{j}^{n}+\left(y_{j+1/2}-y_{j}\right)\frac{\bm{U}_{j+1}^{n}-\bm{U}_{j}^{n}}{y_{j+1}-y_{j}},$ (96)
$K_{j+1/2}^{n}=K_{j}^{n}+\left(y_{j+1/2}-y_{j}\right)\frac{K_{j+1}^{n}-K_{j}^{n}}{y_{j+1}-y_{j}},$
$\epsilon_{j+1/2}^{n}=\epsilon_{j}^{n}+\left(y_{j+1/2}-y_{j}\right)\frac{\epsilon_{j+1}^{n}-\epsilon_{j}^{n}}{y_{j+1}-y_{j}},$
$\left(\frac{\partial\epsilon}{\partial y}\right)^{n}_{j+1/2}=\left(y_{j+1/2}-y_{j}\right)\frac{\epsilon_{j+1}^{n}-\epsilon_{j}^{n}}{y_{j+1}-y_{j}}.$

### v2

@@ 5.3 Non-equilibrium characteristics and velocity distribution functions
The non-equilibrium and non-Newtonian characteristics predicted by the present kinetic model can be essentially explained from the VDFs. We now examine the reduced VDFs obtained from the LR-BGK model at $\text{Re}=10133$, comparing the computed VDFs with their corresponding local equilibrium distributions to quantify the degree of departure from equilibrium. Figure 7 presents the reduced velocity distribution functions $\Phi_{1}$, $\Phi_{2}$, and $\Phi_{3}$ as functions of microscopic velocity $\xi_{y}$ in various normal wall positions ranging from the viscous sublayer to the centerline of the channel, compared against the local equilibrium VDFs derived from the local mean velocity and TKE. Here $\xi_{y}$ in the region $\left[-3\xi_{m},3\xi_{m}\right]$ with $\xi_{m}=\sqrt{4K/3}$ represents the most probable velocity scale based on the local turbulent kinetic energy. The fourth VDF $\Phi_{4}$, which exhibits behaviour similar to $\Phi_{1}$ by remaining close to equilibrium (not shown here for brevity).
@@
The most striking observation is the markedly different behaviours exhibited by the various reduced VDFs. The distribution function $\Phi_{1}$, which represents the zeroth-order moment in the wall-normal direction, remains remarkably close to its local equilibrium state across all wall-normal positions. This near-equilibrium behaviour indicates that the mass conservation constraint is well satisfied throughout the flow domain, consistent with the incompressibility condition. In sharp contrast, $\Phi_{2}$ and $\Phi_{3}$ exhibit pronounced departures from their respective equilibrium distributions, particularly in the outer region of the flow. Here the ”equilibrium” baseline refers to the local isotropic Gaussian distribution $F^{eq}$ (and its reduced counterparts $\Phi^{eq}_{k}$) constructed from the local mean velocity and TKE. Therefore the deviations primarily reflect shear-induced anisotropy and wall-normal inhomogeneity in the averaged turbulence dynamics. The VDF $\Phi_{2}$, associated with the streamwise momentum transport, shows increasing non-equilibrium effects as the distance from the wall increases. Notably, at $y^{+}\approx 419$ and near the channel centerline, $\Phi_{2}$ exhibits a clear departure from the isotropic Gaussian baseline, characterized by an odd-in-$\xi_{y}$ (sign-changing) deviation. This feature is directly related to the Reynolds shear stress, since $\sigma_{xy}=-\langle u_{x}^{\prime}u_{y}^{\prime}\rangle$ is obtained as the moment of the reduced distribution (in our definition, $\sigma_{xy}=-\int\xi_{y}\Phi_{2}d\xi_{y}+U_{x}U_{y}$ for $U_{y}\approx 0$). The observed positive deviation for $\xi_{y}<0$ and negative deviation for $\xi_{y}>0$ contributes to the negative correlation $\langle u_{x}^{\prime}u_{y}^{\prime}\rangle$, consistent with the leading-order (first-order CE) closure and also with linear eddy-viscosity modeling. Similarly, $\Phi_{3}$, which is related to the streamwise kinetic energy transport, demonstrates substantial non-equilibrium characteristics with a comparable spatial evolution pattern. The VDF $\Phi_{3}$ shows progressively stronger departures from equilibrium as the wall-normal distance increases, with the outer region exhibiting the most pronounced non-equilibrium effects. At $y^{+}\approx 419$ and near the channel centerline, $\Phi_{3}$ exhibits a clear asymmetry/shift relative to the isotropic baseline. This feature indicates a nonzero wall-normal transport of TKE flux, which corresponds to a higher-order velocity moment of the VDF and is captured at leading order by the first-order CE closure.
While the leading-order trends of these effects can be understood qualitatively within the first-order CE closure, we further quantify the degree of departure from equilibrium and assess the contributions beyond the first-order closure by examining the second-order moment of $F-F^{eq}$,
@@
Figure 8 presents the deviatoric part of $\bm{\sigma}^{neq}$ predicted by LR-BGK model at $\text{Re}=10133$ and compare them with the CE expansion results given by equation (26). It can be seen that the deviatoric Reynolds stress shows good agreement between the LR-BGK model and the CE expansion throughout the flow domain, validating the theoretical consistency of the kinetic approach. The minor discrepancies observed can be attributed to higher-order terms in the CE expansion that are not captured by the second-order analysis. Furthermore, the deviatoric Reynolds normal stress components better reflect the degree of departure from non-equilibrium of the VDFs mentioned above. In the present model calculations, the deviatoric normal stress components remain close to zero in the viscous sublayer. Across the buffer region, they begin to deviate from zero, with $\sigma^{d}_{xx}$ increasing to positive values while $\sigma^{d}_{yy}$ and $\sigma^{d}_{zz}$ decrease to negative values. Their magnitudes continue to grow toward the outer layer, where the non-equilibrium signatures become more evident. This relatively weak near-wall non-equilibrium signature may also reflect a limited capability of the present formulation to represent buffer-layer non-equilibrium effects, which could contribute to the relatively large discrepancies in the normal stress predictions observed in figure 6 in the near-wall region ($1\lesssim y^{+}\lesssim 10$) compared with DNS data. A possible cause is that, especially in the near-wall region, the relaxation time $\tau$ in the present single-relaxation-time BGK closure is chosen for obtaining the desirable shear stress, as a result, the effective turbulent Knudsen number ($\sim\tau|S|$) becomes too small, so that all other Reynolds stress components are suppressed. One of the future investigations is how to adjust $\tau$ to match shear stress without sacrificing the non-equilibrium effects for normal stress components, including generalization of the single relaxation time collision model to involve multiple relaxation times.
Taken together, the VDFs and their associated non-equilibrium moments provide a framework-specific interpretation for the ability of the present kinetic models to represent anisotropic Reynolds stress behaviour and related non-Newtonian effects. Within this interpretation, departures from local equilibrium reflect finite-relaxation effects in the kinetic closure, analogous in spirit to non-equilibrium transport in rarefied-gas kinetics.
@@
Figure 7: The reduced velocity distribution functions $\Phi_{1}$ (first column), $\Phi_{2}$ (second column) and $\Phi_{3}$ (third column) from the LR-BGK model at $\text{Re}=10133$ as a function of the microscopic velocity $\xi_{y}$ at different wall-normal positions.
@@
Figure 8: Deviatoric Reynolds (a) shear stress component and (b) normal stress components in wall units of turbulent Couette flow at $\text{Re}=10133$. Here, the deviatoric Reynolds stress is defined as $\bm{\sigma}^{d}=\bm{\sigma}-\frac{1}{3}\left(\text{tr}\bm{\sigma}\right)\bm{I}$.
@@ 6 Conclusions
In this work, we extended and analyzed the kinetic theoretic model for turbulence derived from the first principles of incompressible Navier-Stokes equation by chen2023average and demonstrated improved applicability and physical consistency. By redefining the relaxation time as $\tau=K/(7\epsilon)$, the Chapman–Enskog expansion of the kinetic model yielded linear and nonlinear eddy viscosity and gradient diffusion closures with transport coefficients ($C_{\mu}=0.0816,\text{Pr}_{T}=0.7$) that are in quantitative agreement with those of conventional turbulence models. The analysis of our kinetic model reveals that its second-order expansion generates various material derivatives and higher-order derivatives of TKE. To our knowledge, such closed-form Chapman–Enskog expressions for the TKE flux and its nonlinear corrections have not previously been reported for this class of turbulent kinetic models. To address wall-bounded turbulence, we developed two approaches within a unified kinetic framework. For applications outside the viscous sublayer, the non-equilibrium extrapolation boundary condition based on wall function approaches was proposed for the HR-BGK model, maintaining accurate logarithmic layer behaviour at lower computational expense. For direct viscous sublayer resolution, we constructed a low-Reynolds-number kinetic model incorporating a damping function and a viscous diffusion term, with the fully diffuse wall condition.
@@
The present kinetic models were comprehensively validated against experimental data and DNS data for turbulent plane Couette flow, demonstrating accurate prediction of mean velocity profiles, skin friction coefficients, and Reynolds shear stress distributions. They can capture essential flow physics including logarithmic law behaviour and Reynolds stress anisotropy. Although not as pronounced as indicated by DNS, the present kinetic theoretic models are shown to predict the separations between the normal stress components. Nevertheless, accurate prediction of individual normal stress components in canonical wall-bounded flows remains a common challenge. A distinctive advantage of the kinetic approach lies in its robust theoretical foundation, requiring only a single relaxation time, whereas the conventional nonlinear eddy viscosity models necessitate multiple adhocad\ hoc empirical coefficients for closure. These findings suggest that future developments should explore higher-order information about the collision term $\mathcal{C}(F)\equiv-\nabla_{\bm{\xi}}\cdot\langle\bm{a}^{\prime}f^{\prime}\rangle$ to achieve more accurate modelling of stress tensor anisotropy. Along this direction, the nonlinear near-wall anisotropy mechanisms highlighted by 1997Prediction and wallin2000explicit may guide extensions such as an ellipsoidal-statistical BGK (anisotropic Gaussian) equilibrium or a dual-relaxation-time formulation with a few extra parameters to regulate anisotropic relaxation.
@@
Analysis of the reduced velocity distribution functions revealed significant departures from the local Gaussian equilibrium distribution, particularly for streamwise momentum and energy transport components. The non-equilibrium moments obtained from the kinetic model show excellent quantitative agreement with second-order Chapman-Enskog predictions, confirming the theoretical self-consistency of the approach. These non-equilibrium characteristics explain the ability of the kinetic model to capture non-Newtonian effects that linear eddy viscosity models cannot represent, highlighting the fundamental advantage of the kinetic approach over conventional turbulence modelling strategies.
@@
Although this study focused on simple shear flows, the kinetic framework provides a foundation for modelling more complex turbulent phenomena. Future work should address three-dimensional flows, unsteady effects, and extension to compressible turbulence, fully exploiting the theoretical advantages of the kinetic approach. The kinetic theoretic based turbulence model represents a significant step towards describing the fundamental physics of turbulence, involving a wide range of scales without scale separations, it naturally captures higher order nonequilibrium turbulent flow properties as opposed to needing empirical parameters in traditional models. We wish to emphasize that the present work is to propose a kinetic theory based turbulence model, in the form of a Boltzmann-BGK differential equation, for better and more naturally capturing certain fundamental underlying physics of the averaged dynamics of turbulence, instead of a kinetic theory based computational method for numerically solving any turbulence problems.
@@ Appendix A Derivations of the Reynolds stress and turbulent kinetic energy flux via Chapman–Enskog expansion
$Q^{eq}_{\alpha}$ $=\frac{1}{2}\int\left(\xi_{\alpha}-U_{\alpha}\right)\left(\xi_{\beta}-U_{\beta}\right)^{2}F^{eq}d\bm{\xi}=0,$
@@
$\frac{\partial F^{eq}}{\partial U_{\beta}}$ $=\frac{3\left(\xi_{\beta}-U_{\beta}\right)}{2K^{eq}}F^{eq},\quad\frac{\partial F^{eq}}{\partial K^{eq}}=\frac{3}{2K^{eq}}\left(\frac{\left(\xi_{\alpha}-U_{\alpha}\right)^{2}}{2K^{eq}}-1\right)F^{eq},$ (65)
@@
$F^{(1)}=$ $-\tau\left(\partial_{t_{1}}+\xi_{\alpha}\frac{\partial}{\partial x_{1\alpha}}+\bar{a}_{\alpha}^{(1)}\frac{\partial}{\partial\xi_{\alpha}}\right)F^{eq}=-\frac{3\tau\left(\xi_{\alpha}-U_{\alpha}\right)}{2K^{eq}}\left(\frac{\left(\xi_{\alpha}-U_{\alpha}\right)^{2}}{2K^{eq}}-\frac{5}{3}\right)\frac{\partial K^{eq}}{\partial x_{1\alpha}}F^{eq}$ (66)
@@
$\sigma^{(1)}_{\alpha\beta}=-\frac{2}{3}K^{(1)}\delta_{\alpha\beta}+\frac{4\tau K^{eq}}{3}S_{\alpha\beta}^{(1)},\quad Q_{\alpha}^{(1)}=-\frac{10}{9}\tau K^{eq}\frac{\partial{K^{eq}}}{\partial x_{1\alpha}}.$ (67)
@@
$\Phi_{1}\left(y,\xi_{y},t\right)$ $=\int F\left(\bm{x},\bm{\xi},t\right)d\xi_{x}d\xi_{z},\quad\quad\Phi_{2}\left(y,\xi_{y},t\right)=\int\xi_{x}F\left(\bm{x},\bm{\xi},t\right)d\xi_{x}d\xi_{z},$ (74)
$\Phi_{3}\left(y,\xi_{y},t\right)$ $=\int\xi_{x}^{2}F\left(\bm{x},\bm{\xi},t\right)d\xi_{x}d\xi_{z},\quad\Phi_{4}\left(y,\xi_{y},t\right)=\int\xi_{z}^{2}F\left(\bm{x},\bm{\xi},t\right)d\xi_{x}d\xi_{z}.$
@@
$\sigma_{xx}\left(y,t\right)$ $=-\int\left(\Phi_{3}-U_{x}^{2}\Phi_{1}\right)d\xi_{y}+2U_{x}^{2},\quad\sigma_{xy}\left(y,t\right)=-\int\xi_{y}\Phi_{2}d\xi_{y}+U_{x}U_{y},$ (76)
$\sigma_{yy}\left(y,t\right)$ $=-\int\left(\xi_{y}-U_{y}\right)^{2}\Phi_{1}d\xi_{y},\qquad\qquad\sigma_{zz}\left(y,t\right)=-\int\Phi_{4}d\xi_{y}.$
@@
where $\Delta y=y_{j+1/2}-y_{j-1/2}$ is the length of the cell $j$, $\Phi_{j-1/2}^{n+1/2}$ and $\Phi_{j+1/2}^{n+1/2}$ are the left and right VDFs at the cell interface at time $t^{n+1/2}$, respectively. To eliminate the implicit collision term, we introduce the transformed VDFs $\tilde{\Phi}_{\alpha}=\Phi_{\alpha}-\frac{\Delta t}{\tau}\left(\Phi^{eq}_{\alpha}-\Phi_{\alpha}\right)$ and the mean variables can be computed by taking moments of the transformed VDFs,
@@
As a result, the original VDFs and the mean variables at $t^{n+1}$ can be obtained from $\tilde{\Phi}^{n+1}_{\alpha,j}$,
@@
With the Taylor expansion around the cell interface $y_{j+1/2}$, $\Phi_{\alpha}\left(y_{j+1/2},\xi_{y},t^{n+1/2}\right)$ can be approximated by linear interpolations,
@@
Then, $\epsilon$ at the cell interface are also obtained by linear interpolation,
$\epsilon_{j+1/2}^{n}=\epsilon_{j}^{n}+\left(y_{j+1/2}-y_{j}\right)\frac{\epsilon_{j+1}^{n}-\epsilon_{j}^{n}}{y_{j+1}-y_{j}},\quad\left(\frac{\partial\epsilon}{\partial y}\right)^{n}_{j+1/2}=\left(y_{j+1/2}-y_{j}\right)\frac{\epsilon_{j+1}^{n}-\epsilon_{j}^{n}}{y_{j+1}-y_{j}}.$ (88)